OpenStudy (anonymous):

Solve the system using substitution. 2x - y = 0 x + y = 6 What is the solution for x?

OpenStudy (anonymous):

just making sure it was right!

OpenStudy (anonymous):

x=6-y 2(6-y)-y=0 12-y=0 y=12. 2(x)-12=0 2x=12 x=6